OpenStudy (anonymous):

f you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system? Part 2: Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

1) they were the same line.

OpenStudy (anonymous):

1) There are an infinite number of solutions.

OpenStudy (anonymous):

2) \[3y-2x=11\] \[y+2x=9\] add in columns get \[4y=20\] divide get \[y=5\] substitute 5 for y in equation 2 get \[5+2x=9\] \[2x=4\] \[x=2\]

OpenStudy (anonymous):

do you know the 3 cases for solving systems? infinite, no solution, and one solution? that will be on the exam

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

well you see where the lines are the same and they overlap that is something like 9=9 or -10=-10 thats infinite

OpenStudy (anonymous):

if you get 0=7 or 3=-25 then no solution